If the rate is at 1.25% effective, how long will Php 5,000 become Php 5,500?





1
Expert's answer
2022-01-12T18:14:56-0500

find the nominal rate

"R=(1+r)^n-1"

"1.25=(1+r)^1-1"

"2.25=1+r"

r=1.25Since the number of years of the deposit is a degree, then it will find the term of increase through the logarithm using the formula

"t=\frac{ln N}{ln(1+r)}"


"N=\frac{5500}{5000}=1.1"


r=0.0125


"t=\frac{ln N}{ln(1+r)}=\frac{ln 1.1}{ln(1+0.0125)}=\frac{0.095310179804325}{0.012422519998557}=7.9"